OpenStudy (anonymous):

find all cube roots of -8 answer in a+bi form

OpenStudy (phi):

In polar form \[ -8 = 8e^{i\pi} \] and \[ -8^{\frac{1}{3}}= 8^{\frac{1}{3}}e^{\frac{i\pi}{3}} = 2e^{\frac{i\pi}{3}} \] Use \[ Ae^{iB}= Acos(B)+iAsin(B) \] to change to rectangular form. To get the other roots (for a total of 3), multiply -8 by 1 as in \( 8e^{i\pi} \cdot e^{i2\pi} \) and then raise to the 1/3. Each time you multiply by e^i2pi you will get another root, until you return to your first value.